Shift report 20011118 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Nov 18 2001 - 15:55:05 EST

  • Next message: Apache: "Shift report 20011119 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Nov 18 15:55:04 2001, Shift: 08:00-16:00
    Supervisor: J.I.Jordre, with: 
    Log book#11 pages 8 - 8
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from 8:12, but steering until 9:15. Store ended 12:25 after
    controlled dump. MCR said they would do injection studies 12:30-13:30, 
    but studies continued until ~14:30. Quench link interloch and FEC
    controls problems after that. No estimate for beam yet.
    2. MCR @ 90 deg. 6 B, FS @ 12 deg. 2/3 A:
       Run 5873	 92k tr. 6	(taken during steering)
       Run 5874	196k tr. 6
       Pedestal, DVM and TPC pedestal runs 5875, 5876, 5878
    3. No new detector problems. Had to restart RunServer and IOServer
    once when RunServer reported that Camac and FB4 were not connected. No
    problem after that.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Nov 18 2001 - 15:55:46 EST